A manufactured product has the following information for June. Standard Actual Direct materials (7 lbs. @ $9 per lb.) 56,500 lbs. @ $9.10 per lb. Direct labor (3 hrs. @ $16 per hr.) 23,700 hrs. @ $16.50 per hr. Overhead (3 hrs. @ $11 per hr.) $ 270,100 Units manufactured 8,000

Compute the direct labor rate variance and the direct labor efficiency variance. Indicate whether each variance is favorable or unfavorable. AH = Actual Hours SH = Standard Hours AR = Actual Rate SR = Standard Rate

Actual Cost

Standard Cost
Answer #1

Direct labor rate variance = (SR-AR)*AH

= (16-16.50)*23700 = $11850 Unfavorable

Direct labor efficiency variance = (SH - AH)*SR

= (8000*3 - 23700)*16 = $4800 Favorable